OpenStudy (anonymous):

What is -(-5)-5(9-8) plzz explain to me step by step

OpenStudy (appleduardo):

remember: - & - = + - & + =- + & - = - + & + =+

OpenStudy (appleduardo):

so --5 = 5 -5 = 0 * (9-8) =0

OpenStudy (anonymous):

Wait. What????

OpenStudy (appleduardo):

you have a '-' before "-5" so: -(-5) = 5, then: 5-5 = 0 ... then make the 9-8, that gives us: 1, but 1 times 0 (which was the result of: -(-5)-5) equals to 0

OpenStudy (anonymous):

Ohh ok. Ok. So... You make the -(-5) turn into a positive

OpenStudy (appleduardo):

yeep!

OpenStudy (anonymous):

Ohh ok Thx